I have three questions:
1.The amount of money in a savings account decreases from $340 to $220. Find the percent decrease rounded to the nearest tenth of a percent.
2.Acme Manufacturing's daily production of widgets is 140% of its competition's. If the competition produces 1230 units per day, how much does Acme produce?
3. Grace has 85% of total points possible in her science class. If there are 350 points possible, how many points does Grace have?

Answer by Theo(13342)

1.The amount of money in a savings account decreases from $340 to $220. Find the percent decrease rounded to the nearest tenth of a percent.

340 - 220 = 120
the amount of the decrease is 120.
the decrease is from the original amount of 340.
the percent decrease is 120 / 340 = .3529411765 * 100 = 35.29411765% which is rounded to 35.3%.
you start with 340.
multiply that by .3529411765 to get 120.
subtract 120 from 340 to get 220.
the percent decrease is 35.3% rounded to the nearest tenth of a percent.

2.Acme Manufacturing's daily production of widgets is 140% of its competition's. If the competition produces 1230 units per day, how much does Acme produce?

140% of 1230 is equal to 140/100 * 1230 which is equal to 1.4 * 1230 which is equal to 1722.
1722/1230 = 1.4 * 100 = 140%.

3. Grace has 85% of total points possible in her science class. If there are 350 points possible, how many points does Grace have?

85% of 350 is equal to 85/100 * 350 which is equal to .85 * 350 which is equal to 297.5
grace has 297.5 points out of 350.
297.5/350 = .85 * 100 = 85%